Two water taps can togetheir fill a tank in 75/8 hours.the tap with larger diameter takes 10 hour less than the smaller one separately to fill the tank.find the time taken by the both tap to fill a tank separately
Answer by stanbon(75887)

Two water taps can togetheir fill a tank in 75/8 hours.the tap with larger diameter takes 10 hour less than the smaller one separately to fill the tank.find the time taken by the both tap to fill a tank separately
--------------
together DATA:: time = 75/8 hrs/job ; rate = 8/75 job/hr
------
smaller tap DATA: 8time = x hrs/job ; rate = 1/x job/hr
------
larger tap data: time = x-10 hrs/job ; rate = 1/(x-10) job/hr
-------
Equation:
rate + rate + together rate
1/x + 1/(x-10) = 8/75
----
75(x-10) + 75x = 8x(x-10)
150x + 750 = 8x^2 - 80x
----
8x^2 - 230x - 750 = 0
------
x = 31.7 hrs (time for smaller tap to do the job alone)
x-10 = 21.7 hrs (time for larger tap to do the job alone)
------
cheers,
Stan H.
==========